OpenStudy (anonymous):

A 70.0 kg sailor climbs a 28.3 m rope ladder at an angle of 45.0° with the mast. How much work did he do? A. 19,400 J B. 686 J C. 1,400 J D. 13,700 J :/

OpenStudy (michele_laino):

the requested work W is: \[W = mgd\cos \left( {45} \right) = 70 \times 9.81 \times 28.3 \times \cos 45\]

OpenStudy (michele_laino):

since we havfe to consider the projection of the weight force, along the direction of the displacement

OpenStudy (michele_laino):

have*

OpenStudy (anonymous):

ohh okay! and so we calculate and get 13741? so our solution is D? 13,700 J since it is approximation?

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

yay!! thank you!!

OpenStudy (michele_laino):

ok! :)